A rectangular parcel of land is 50 ft wide. The length of a diagonla between opposite corners is 10 ft more than the lenght of the parcel. What is the lenght of the parcel?
Answer by ilana(307)

If we draw the rectangle with width=50, length=x, diagonal=x+10, we find a right triangle with sides 50 and x, and hypotenuse=x+10. We then use the Pythagorean Theorem to get (x+10)^2 = 50^2 + x^2, so x^2+20x+100=2500+x^2.
When we subtract x^2 from both sides, we get 20x+100=2500, subtracting 100 gives 20x=2400, dividing by 20 gives x=120ft. So this is the length.